40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6597     Entry time: Thu May 3 13:35:56 2012
Author: Jenne 
Type: Bureaucracy 
Category: General 
Subject: 40m Meeting Action Items 

Action Items:

IFO:

         Arm cavity sweeps, mode scan - JENNE

        Align AS OSA (others?) - JENNE       

        Investigate PRMI glitches, instability (take PRM oplev spectra locked, unlocked, to see if PRM is really moving) - KOJI, JENNE, DEN

         Connect up beatbox for Xarm use - KOJI with JENNE       

 

TT:

Prepare electronics for TTs (coil drivers) - JAMIE

In-air TT testing to confirm we can control / move TTs before we vent (starting in ~2 weeks) - SURESH

Connect TTs to digital system and controls, lay cables if needed - JAMIE with SURESH

OAF:

OAF comparison plot, both online and offline, comparing static, adaptive and static+adaptive - DEN

Static-only OAF noise budget (Adaptive noise budget as next step) - DEN

Build amplifiers for new small microphones - DEN 

SLC:

Black glass: to clean&bake - KOJI

Scattered light measurement at the end stations: design / confirmation of the mechanical parts/optics/cameras - JAN

MM:

IPPOS beam measurement - SURESH with JENNE

AS beam measurement (if beam is bright enough) - SURESH and JENNE

Mode matching calculations, sensitivity to MC waist measurement errors, PRM position - JENNE

Think up diagnostic measurement to determine mode matching to PRC while chambers are open, while we tweak MMT - JAMIE, JENNE, KOJI, SURESH

Computers:

       Upgrade Rossa, Allegra to Ubuntu, make sure Dataviewer and DTT work - JAMIE

 

ELOG V3.1.3-